# BanditSpec: Adaptive Speculative Decoding via Bandit Algorithms

### v1

@@ G.2 Experiments on Larger Models
In addition to the two models in the main paper, we conduct an addtional set of experiment on a larger target model, namely LLaMA-2-13B (Touvron et al., 2023). As Table 1 indicates Eagle-2 (Li et al., 2024b) is one of the best speculative decoding methods, we adopt it as the baseline. The other setups are the same as the ones in Section 6.1.
@@
Table 3: Empirical Comparison between the proposed algorithms and Eagle-2 (Li et al., 2024b) with LLaMA-2-13B as the target model, measured by Mean Accepted Tokens (MAT) ($\uparrow$) and Tokens/s ($\uparrow$). The best result is highlighted in bold, while the second best result is underlined. The proposed algorithms remain effective on larger models.
@@ G.3 Experiments on Different Hardwares
In the main paper, the experiments are conducted on a single A100 GPU. In this section, we conduct an additional set of experiment on GeForce RTX 4090. We adopt Eagle-2 (Li et al., 2024b) as the baseline and Spec Bench (Xia et al., 2024) as the benchmark. The result is presented in Table 4. We observe a similar trend as the result presented in Table 1. The proposed method remains useful with a different hardware setup.

### v2

@@ G.2 Experiments on Larger Models
In addition to the two models in the main paper, we conduct an additional set of experiments on a larger target model, namely LLaMA-2-13B (Touvron2023Llama2O). As Table 1 indicates, Eagle-2 (li2024eagle2) is one of the best speculative decoding methods, we adopt it as the baseline. The other setups are the same as the ones in Section 6.1.
@@
Table 3: Empirical Comparison between the proposed algorithms and Eagle-2 (li2024eagle2) with LLaMA-2-13B as the target model, measured by Mean Accepted Tokens (MAT) ($\uparrow$) and Tokens/s ($\uparrow$). The best result is highlighted in bold, while the second best result is underlined. The proposed algorithms remain effective on larger models.
@@ G.3 Experiments on Different Hardwares
In the main paper, the experiments are conducted on a single A100 GPU. In this section, we conduct an additional set of experiments on GeForce RTX 4090. We adopt Eagle-2 (li2024eagle2) as the baseline and Spec Bench (xia2024unlocking) as the benchmark. The result is presented in Table 4. We observe a similar trend as the result presented in Table 1. The proposed method remains useful with a different hardware setup.